# Ephemerality meets LiDAR-based Lifelong Mapping

### v1

@@ I INTRODUCTION
Over the past decade, LiDAR (LiDAR)-based mapping has significantly advanced [1, 2, 3, 4], increasing the demand for long-term deployment of such systems in various fields including urban areas or construction sites [5]. These environments are inherently dynamic; objects frequently move, and layouts change. To handle these dynamics, continuously revisiting and maintaining the map of the environment—lifelong mapping—is required.
LiDAR-based lifelong mapping has gained interest relatively recently compared to the visual domain [6, 7, 8]. Long-term mapping pipeline for static map [9] or semantic map [10] construction were suggested, but the standard framework for lifelong mapping was absent. Recently, the modular lifelong mapping frameworks [11, 12] were suggested. They process the session—a set of point clouds and poses—as an input and focus on the inter-session changes for efficient map management and incremental update.
@@ II-A Change Detection
II-B Lifelong Mapping
@@ III-A System Overview
Figure 2: Overview of the ELite system pipeline. Given multiple input sessions, ELite updates the map by estimating local ephemerality within each session and updating global ephemerality across sessions. The system operates through three modules: multi-session map alignment, dynamic object removal, and map update. Additionally, the system manages three types of maps: lifelong map, delta map, and static map.
ELite manages two stages of ephemerality: local ephemerality ($\epsilon_{l}$) and global ephemerality ($\epsilon_{g}$). $\epsilon_{l}$ reflects the probability of a point being dynamic within a session (e.g., moving cars exhibit high $\epsilon_{l}$ than the parked car), while $\epsilon_{g}$ captures the long-term likelihood of a point being transient (e.g., temporally parked car exhibit high $\epsilon_{g}$ than the building).
Fig. 2 illustrates an overview of our system. Starting from the base map $\mathcal{M}_{1}$ constructed directly from the first session $\mathcal{S}_{1}$, our lifelong mapper $L(\cdot)$ incrementally updates previous lifelong map $\mathcal{M}_{t-1}$ with new session $\mathcal{S}_{t}$.
@@
Lifelong map refers to the map point cloud where each point contains $(x,y,z,\epsilon_{g})$, and session $\mathcal{S}=\left\{\left(\mathcal{P}_{i},\mathbf{T}_{i}\right)\right\}_{i=1}^{N}$ refers to a set of scans and corresponding poses in a local coordinate frame, acquired from an arbitrary positioning system. For example, it may consist of the raw scan of keyframes and corresponding poses from LiDAR odometry [4] or SLAM (SLAM) [3].
ELite first aligns the new session to the existing lifelong map by finding the scan-wise optimal transformation (§III-B). The following dynamic object removal module estimates the local ephemerality $\epsilon_{l}$ for each point using point-wise ephemerality propagation and employs them for removing dynamic points (§III-C). Leveraging the local ephemerality, the map update module classifies map points into several categories and applies category-wise update rule to calculate updated global ephemerality $\epsilon_{g}$-D). After a sufficient number of updates, $\epsilon_{g}$ can act as a weight for aligning new sessions into the lifelong map, improving the robustness of the whole system.
@@ III-B Multi-session Map Alignment
Figure 3: Illustration of our map alignment module, which starts by aligning poses using the initial transform from the first loop closure candidate. It then refines alignment in two stages: forward and backward. By iterating through scans in forward and reverse order, the module updates poses using scan-to-map ICP registration, ensuring global and local consistency in the final pose estimates.
The goal here is to find the optimal transformation $\mathbf{T}^{\prime}_{i}$ for each scan $\mathcal{P}_{i}$ in a session and pass the aligned session $\mathcal{S^{\prime}}=\left\{\left(\mathcal{P}_{i},\mathbf{T}^{\prime}_{i}\right)\right\}_{i=1}^{N}$ to the subsequent modules.
An overview of our alignment module is illustrated in Fig. 3. To begin, we select a loop pair $(\mathcal{P}^{\text{sess}}_{s},\mathcal{P}^{\text{map}}_{m})$ using Scan Context [39] and estimate initial transformation $\mathbf{T}^{\text{init}}$. By iterating the indices $i\in[s,N]$, $\mathbf{T}^{\text{init}}$ is refined with the transformation $\mathbf{T}_{i}^{\text{ICP}}$ calculated with GICP (GICP) [40]. $\mathbf{T}_{i}^{\text{ICP}}$ is then applied to point clouds from $i$th to the end index $N$, providing better initial guesses for the subsequent ones. This step is called the forward registration, as the transformation propagates forward through the point clouds. The refined pose for each point cloud is:
$\vspace{-1mm}\mathbf{T}^{\text{fwd}}_{i}=\prod_{j=s}^{i}\mathbf{T}_{j}^{\text{ICP}}\cdot\mathbf{T}^{\text{init}}\cdot\mathbf{T}_{i}\quad\text{if}\ i\in[s,N]\vspace{-1mm}$ (2)
However, the point clouds before index $s$ may still be misaligned, and there may exist accumulated errors during the forward registration. To address this, we apply a backward registration, where point clouds are registered in reverse order. The final pose after the backward process is:
$\mathbf{T}^{\prime}_{i}=\prod_{k=i}^{N}\mathbf{T}_{k}^{\text{ICP,rev}}\cdot\prod_{j=s}^{i}\mathbf{T}_{j}^{\text{ICP}}\cdot\mathbf{T}^{\text{init}}\cdot\mathbf{T}_{i}\vspace{-1mm}$ (3)
This two-step alignment process, moving forward and then backward, is similar to “zipping up” two maps. By using both directions, the system achieves consistent alignment, even if there is drift in the trajectory of the new session.
During scan-to-map registration, static map points will exhibit low $\epsilon_{g}$ values and be more reliable for registration, while the ephemeral points will retain higher $\epsilon_{g}$ values. The alignment module uses $\epsilon_{g}$ as the weight in the GICP optimization:
$\mathbf{T}^{\text{ICP}}=\underset{\mathbf{T}}{\mathrm{argmin}}\sum\limits_{i}\left(1-\epsilon_{g,i}\right)\lVert\mathbf{d}_{i}(\mathbf{T})\rVert^{2}_{\mathbf{\Sigma}_{i}}$ (4)
The details of $\epsilon_{g}$, including its derivation within our framework, are provided in §III-C and §III-D. As more sessions are used for updates, the lifelong map will exhibit a more reliable $\epsilon_{g}$ distribution. This, in turn, steadily improves the robustness of our system during long-term operations.
@@ III-C Dynamic Object Removal
Figure 4: As shown within the green square, voxelization discretizes the space and inflates the occupied area, leading to potential errors in classifying dynamic points. In comparison, our method can be more precise in removing dynamic objects by directly utilizing rays and updating point-wise ephemerality.
Dynamic object removal module first aggregates the scans to construct session map $\mathcal{M}_{\mathcal{S}^{\prime}_{t}}=\bigcup_{i=1}^{N}\mathbf{T}^{\prime}_{i}\mathcal{P}_{i}$. It aims to recover cleaned session map ^$\hat{\mathcal{M}}_{S^{\prime}_{t}}$ by removing dynamic points from ${\mathcal{M}}_{S^{\prime}_{t}}$. Viewing a scan as a bundle of rays, we iteratively update the local ephemerality $\epsilon_{l}$ of each point in $\mathcal{M}_{\mathcal{S^{\prime}}_{t}}$ using these rays. Local ephemerality of the point should decrease if the point is close to the occupied space—endpoint of each ray, and increase if located in the free space—path along the rays.
Endpoints in each scan $\mathcal{P}_{i}$ is gathered into observed points $\mathcal{O}_{i}$ and free points $\mathcal{F}_{i}$ are sampled from the rays to approximate the free space [41]. We assume that $\epsilon_{l}$ around each observed point $\mathbf{o}_{i}$ or free point $\mathbf{f}_{i}$ follows one-dimensional Gaussian distribution with respect to the distance $x$:
$f(x)=\left\{\begin{array}[]{lll}\min\left(\alpha\cdot\left(1-\exp\left\{-\dfrac{x^{2}}{{\sigma_{\mathbf{o}}}^{2}}\right\}\right)+\beta,\alpha\right)&\text{for}&\mathbf{o}_{i}\in\mathcal{O}_{i}\\ \max\left(\alpha\cdot\left(1+\exp\left\{-\dfrac{x^{2}}{{\sigma_{\mathbf{f}}}^{2}}\right\}\right)-\beta,\alpha\right)&\text{for}&\mathbf{f}_{i}\in\mathcal{F}_{i}\\ \end{array}\right.$
where $\alpha$ and $\beta$ are scale parameters (fixed to 0.5 and 0.1, respectively) and $\sigma$ is standard deviation. Starting from the initial value $\epsilon_{l}=0.5$, each point in $\mathcal{O}_{i}$ or $\mathcal{F}_{i}$ retrieves its k-nearest neighbors (kNN) points in $\mathcal{M}_{\mathcal{S^{\prime}}_{t}}$ and updates their $\epsilon_{l}$ through Bayesian inference:
$\centering\begin{split}\epsilon_{{l},\text{new}}=\frac{f(x)\cdot\epsilon_{{l},\text{prev}}}{f(x)\cdot\epsilon_{{l},\text{prev}}+(1-f(x))\cdot(1-\epsilon_{{l},\text{prev}})}\end{split}\@add@centering$ (6)
As illustrated in Fig. 4, directly utilizing the ray relieves the need for discretizing the space and allows more precise propagation of ephemerality than its voxel-based counterparts. After a series of updates, points in $\mathcal{M}_{\mathcal{S}^{\prime}_{t}}$ with $\epsilon_{l}>\tau_{l}$ are regarded as dynamic points thus removed, yielding cleaned session map ^$\hat{\mathcal{M}}_{S^{\prime}_{t}}$.
@@ III-D Long-term Map Update
The map update module first aggregates the previous lifelong map and current cleaned session map into the lifelong map $\mathcal{M}_{t}=\mathcal{M}_{t-1}\cup\hat{\mathcal{M}_{S^{\prime}_{t}}}$ and classifies the points into five categories using nearest neighbors (NNs) search, as shown in Fig. 5: (i) Coexisting points $\mathcal{C}_{t}$, (ii) Deleted points $\mathcal{D}_{t}$ in $\mathcal{M}_{t-1}$, (iii) Emerged points $\mathcal{E}_{t}$ in ^$\hat{\mathcal{M}}_{S^{\prime}_{t}}$. Points in either $\mathcal{D}_{t}$ or $\mathcal{E}_{t}$ are further classified into: (iv) Previously explored points, and (v) Newly explored points, if they are located in areas not explored by the previous and current sessions.
@@
Each point in $\mathcal{C}_{t}$ updates its $\epsilon_{g}$ through Bayesian inference using the previous $\epsilon_{g}$ of $\mathcal{M}_{t-1}$ and the current $\epsilon_{l}$ in ^$\hat{\mathcal{M}}_{S^{\prime}_{t}}$:
$\centering\epsilon_{{g_{t}}}=\frac{\epsilon_{{g_{t-1}}}\cdot\epsilon_{l_{t}}}{\epsilon_{{g_{t-1}}}\cdot\epsilon_{l_{t}}+(1-\epsilon_{{g_{t-1}}})\cdot(1-\epsilon_{l_{t}})}\@add@centering$ (7)
$\mathcal{D}_{t}$ not only include the disappeared objects but also erroneous points due to sensor noise, pose error, etc. In order to focus on meaningful changes, we introduce an objectness factor $\gamma$, which can prioritize the object-wise changes and suppress the others. We empirically find that the density $\rho$ of the points, approximated by normalizing the number of neighbors within radius $r$, can be effective for discerning the object-wise changes and the others. We therefore choose $\gamma=\rho^{\frac{1}{3}}$ and use it to increase $\epsilon_{g}$ of each point in $\mathcal{D}_{t}$ through Bayesian update, similar to (7).
Points in $\mathcal{E}_{t}$ incorporate various cases, such as newly constructed buildings or temporarily parked cars, introducing uncertainty about their permanence. To address this, we scale their local ephemeralities by an uncertainty factor $k$, along with the objectness factor $\gamma$:
@@
$\epsilon_{{g,i}}=k\cdot\left(2-\gamma\right)\cdot\epsilon_{{l,i}},\quad\forall i\in\mathcal{E}_{t}$ (8)
@@
Points in category (iv) do not update their previous $\epsilon_{g}$, and points in (v) initialize $\epsilon_{g}$ with their $\epsilon_{l}$ value.
As more observations accumulate, the $\epsilon_{g}$ of new static points will decrease and stabilize compared to that of ephemeral objects. Updated points in $\mathcal{M}_{t}$ with their corresponding changes ($\Delta\epsilon_{g}$) are logged into delta map of session $\mathcal{S}_{t}$. Delta map differs from the diff map suggested in LT-mapper [11] in that it continuously tracks and records the amount of changes. This precisely encodes subtle differences between the environment changes and can be further utilized for object-wise spatial analysis. After the update, the user can query a static map by retaining elements in $\mathcal{M}_{t}$ if their $\epsilon_{g}$ value is below the user-adjustable threshold $\tau_{g}$.
@@
Figure 5: In our map update module, points are classified into five categories: coexisting points ($\mathcal{C}_{t}$) are shown in grey; deleted points ($\mathcal{D}_{t}$) are red if truly removed and pink if from previous-only visited regions; emerged points ($\mathcal{E}_{t}$) are blue if newly added and sky blue if observed only in currently visited regions. Each category follows a distinct update strategy for effective map updates.

### v2

@@ I INTRODUCTION
Over the past decade, LiDAR (LiDAR)-based mapping has significantly advanced [1, 2, 3, 4], increasing the demand for long-term deployment of such systems in various fields, including urban areas or construction sites [5]. These environments are inherently dynamic; objects frequently move, and layouts change. To handle these dynamics, continuously revisiting and maintaining the map of the environment—lifelong mapping—is required.
LiDAR-based lifelong mapping has gained interest relatively recently compared to the visual domain [6, 7, 8]. Long-term mapping pipelines for static map [9] or semantic map [10] construction were suggested, but the standard framework for lifelong mapping was absent. Recently, the modular lifelong mapping frameworks [11, 12] were suggested. They process the session—a set of point clouds and poses—as an input and focus on the inter-session changes for efficient map management and incremental update.
@@ II-A Change Detection
II-B LiDAR-based Lifelong Mapping
@@ III-A System Overview
Figure 2: Overview of the ELite system pipeline. Given multiple input sessions, ELite updates the map by estimating local ephemerality within each session and updating global ephemerality across sessions. The system operates through three modules: multi-session map alignment, dynamic object removal, and map update. Additionally, the system manages three types of maps: a lifelong map, a delta map, and a static map.
ELite manages two stages of ephemerality: local ephemerality ($\epsilon_{l}$) and global ephemerality ($\epsilon_{g}$). Here, $\epsilon_{l}$ reflects the probability of a point being dynamic within a single session (e.g., moving cars have higher $\epsilon_{l}$ than parked cars), while $\epsilon_{g}$ captures the long-term likelihood of a point being transient (e.g., a car repeatedly parked in the same location exhibits a higher $\epsilon_{g}$ than a permanent building).
Fig. 2 provides an overview of the system. Starting from the base map $\mathcal{M}_{1}$, which is built directly from the first session $\mathcal{S}_{1}$ (see §III-C), our lifelong mapper $L(\cdot)$ incrementally updates the previous lifelong map $\mathcal{M}_{t-1}$ using the new session $\mathcal{S}_{t}$:
@@
The lifelong map is a point cloud in which each point contains $(x,y,z,\epsilon_{g})$. A session $\mathcal{S}=\left\{\left(\mathcal{P}_{i},\mathbf{T}_{i}\right)\right\}_{i=1}^{N}$ is a set of scans and their corresponding poses in a local coordinate frame, obtained from any positioning system (e.g. LiDAR odometry [4] or SLAM [3]).
First, ELite aligns the new session to the existing lifelong map by finding the scan-wise optimal transformations (§III-B). Next, it estimates local ephemerality $\epsilon_{l}$ for each point via point-wise ephemerality propagation in the dynamic object removal module and then discards dynamic points (§III-C). Finally, leveraging the estimated local ephemerality, the map update module classifies map points into multiple categories and applies category-specific update rules to compute the new global ephemerality $\epsilon_{g}$-D). Over repeated updates, $\epsilon_{g}$ acts as a reliable weight for aligning newly acquired sessions, thereby enhancing system robustness during extended operation.
@@ III-B Multi-session Map Alignment
Figure 3: Illustration of our map alignment module, which begins by aligning poses using the initial transform from the first loop closure candidate. It then refines alignment in two stages: forward and backward. By iterating through scans in both directions, the module updates poses via scan-to-map ICP registration, ensuring global and local consistency in the final pose estimates.
The goal of an alignment module is to find the optimal transformation $\mathbf{T}^{\prime}_{i}$ for each scan $\mathcal{P}_{i}$ in a session and pass the aligned session $\mathcal{S^{\prime}}=\left\{\left(\mathcal{P}_{i},\mathbf{T}^{\prime}_{i}\right)\right\}_{i=1}^{N}$ to subsequent modules.
Fig. 3 illustrates our alignment pipeline. We first select a loop pair $(\mathcal{P}^{t-1}_{m},\mathcal{P}^{t}_{s})$ from two sessions ($\mathcal{S}_{t-1}$ and $\mathcal{S}_{t}$) via Scan Context [39] to estimate an initial transformation $\mathbf{T}^{\text{init}}$. We then refine $\mathbf{T}^{\text{init}}$ in a forward pass over indices $i\in[s,N]$ using GICP (GICP) [40]:
$\vspace{-1mm}\mathbf{T}^{\text{fwd}}_{i}=\prod_{j=s}^{i}\mathbf{T}_{j}^{\text{ICP}}\cdot\mathbf{T}^{\text{init}}\cdot\mathbf{T}_{i}\quad\text{if}\ i\in[s,N].\vspace{-1mm}$ (2)
After each GICP step at $i$, $\mathbf{T}_{i}^{\text{ICP}}$ is also applied to subsequent point clouds to provide better initial guesses. Since point clouds prior to index $s$ may remain misaligned and there could be accumulated errors, we perform a backward pass from $i=N$ down to $s$. This yields:
$\mathbf{T}^{\prime}_{i}=\prod_{k=i}^{N}\mathbf{T}_{k}^{\text{ICP,rev}}\cdot\prod_{j=s}^{i}\mathbf{T}_{j}^{\text{ICP}}\cdot\mathbf{T}^{\text{init}}\cdot\mathbf{T}_{i}.\vspace{-1mm}$ (3)
Working in both directions—similar to “zipping up” two maps—improves overall alignment consistency and compensates the trajectory drift.
During scan-to-map registration, points with low $\epsilon_{g}$ (permanent) have higher weight in GICP, while points with high $\epsilon_{g}$ (ephemeral) have less influence.
$\mathbf{T}^{\text{ICP}}=\underset{\mathbf{T}}{\mathrm{argmin}}\sum\limits_{i}\left(1-\epsilon_{g,i}\right)\lVert\mathbf{d}_{i}(\mathbf{T})\rVert^{2}_{\mathbf{\Sigma}_{i}},$ (4)
$\epsilon_{g,i}$ is the global ephemerality of the $i$th map point, and $\mathbf{d}_{i}(\mathbf{T})$ is the residual error under transformation $\mathbf{T}$. Detailed explanation of $\epsilon_{g}$ will be provided in §III-C and §III-D. As the system incorporates more sessions, the $\epsilon_{g}$ distribution in the lifelong map becomes more reliable, steadily improving alignment robustness.
@@ III-C Dynamic Object Removal
Figure 4: When voxelizing space, as indicated by the green square, the occupied area is inflated and errors can occur when a single voxel mixes static and dynamic points. Our method updates point-wise ephemerality based on ray information, enabling more precise removal of dynamic objects.
The dynamic object removal module first aggregates all scans in the aligned session to form $\mathcal{M}_{\mathcal{S}^{\prime}_{t}}=\cup_{i=1}^{N}\mathbf{T}^{\prime}_{i}\mathcal{P}_{i}$. Its goal is to produce a cleaned session map ^$\hat{\mathcal{M}}_{S^{\prime}_{t}}$ by discarding dynamic points from $\mathcal{M}_{\mathcal{S}^{\prime}_{t}}$.
Treating each scan as a set of rays, we iteratively update the local ephemerality $\epsilon_{l}$ of points in $\mathcal{M}_{\mathcal{S}^{\prime}_{t}}$. A local ephemerality of a point should decrease if it is near an occupied space (the endpoint of a ray) and increase if it lies in free space (the path along a ray). We propagate ephemerality across space by modeling a function of the distance $x$ from a ray:
$f(x)=\left\{\begin{array}[]{lll}\min\Bigl{(}\alpha\cdot\bigl{(}1-\exp\{-x^{2}/{\sigma_{\mathbf{o}}}^{2}\}\bigr{)}+\beta,\ \alpha\Bigr{)}&\text{if}&\mathbf{o}_{i}\in\mathcal{O}_{i}\\[7.0pt] \max\Bigl{(}\alpha\cdot\bigl{(}1+\exp\{-x^{2}/{\sigma_{\mathbf{f}}}^{2}\}\bigr{)}-\beta,\ \alpha\Bigr{)}&\text{if}&\mathbf{f}_{i}\in\mathcal{F}_{i}\\ \end{array}\right.$
$\mathcal{O}_{i}$ is the set of endpoints in scan $\mathcal{P}_{i}$ and $\mathcal{F}_{i}$ consists of sampled points along the rays that approximate free space [41]. $\alpha$ and $\beta$ are scale parameters (fixed to 0.5 and 0.1, respectively), while $\sigma_{\mathbf{o}}$ and $\sigma_{\mathbf{f}}$ are standard deviations.
Starting from an initial value $\epsilon_{l}=0.5$, each point in $\mathcal{O}_{i}$ or $\mathcal{F}_{i}$ retrieves its k-nearest neighbors in $\mathcal{M}_{\mathcal{S}^{\prime}_{t}}$ and updates their $\epsilon_{l}$ via Bayesian inference:
$\epsilon_{\mathrm{l,new}}=\frac{f(x)\cdot\epsilon_{\mathrm{l,prev}}}{f(x)\cdot\epsilon_{\mathrm{l,prev}}+(1-f(x))\cdot(1-\epsilon_{\mathrm{l,prev}})}.$ (6)
As shown in Fig. 4, leveraging the actual rays rather than voxelization prevents the artificial inflation of occupied regions and leads to more precise ephemerality propagation. After several updates, points in $\mathcal{M}_{\mathcal{S}^{\prime}_{t}}$ with $\epsilon_{l}<\tau_{l}$ are deemed static and pushed into ^$\hat{\mathcal{M}}_{S^{\prime}_{t}}$.
@@ III-D Long-term Map Update
The map update module merges the previous lifelong map $\mathcal{M}_{t-1}$ with the cleaned session map ^$\hat{\mathcal{M}}_{S^{\prime}_{t}}$ to form the new lifelong map $\mathcal{M}_{t}$:
$\mathcal{M}_{t}=\mathcal{M}_{t-1}\cup\hat{\mathcal{M}}_{S^{\prime}_{t}}.$
@@
It then classifies the points into five categories based on their nearest neighbors (NNs), as shown in Fig. 5: (i) Coexisting points $\mathcal{C}_{t}$, (ii) Deleted points $\mathcal{D}_{t}$ from $\mathcal{M}_{t-1}$, (iii) Emerged points $\mathcal{E}_{t}$ in ^$\hat{\mathcal{M}}_{S^{\prime}_{t}}$. Points in either $\mathcal{D}_{t}$ or $\mathcal{E}_{t}$ are further classified into: (iv) Previously explored points, and (v) Newly explored points, if they are located in areas not explored by both the previous and current sessions.
@@
For each point in $\mathcal{C}_{t}$, we update its $\epsilon_{g}$ via Bayesian inference using the previous $\epsilon_{g}$ (from $\mathcal{M}_{t-1}$) and the current $\epsilon_{l}$ (from ^$\hat{\mathcal{M}}_{S^{\prime}_{t}}$):
$\epsilon_{\mathrm{g_{t}}}=\frac{\epsilon_{\mathrm{g_{t-1}}}\cdot\epsilon_{l_{t}}}{\epsilon_{\mathrm{g_{t-1}}}\cdot\epsilon_{l_{t}}+(1-\epsilon_{\mathrm{g_{t-1}}})\cdot(1-\epsilon_{l_{t}})}.$ (7)
The set $\mathcal{D}_{t}$ can include truly removed objects as well as spurious points from sensor noise or pose errors. To emphasize meaningful changes and suppress noise, we introduce an objectness factor $\gamma$, which prioritizes object-like clusters. We find that the local density $\rho$ of points effectively distinguishes object-level changes from noise and leverage it to define $\gamma$.
$\gamma_{i}=\rho_{i}^{1/3},\quad\rho_{i}\propto\{\mathbf{p}_{j}\in\mathcal{D}\mid\|\mathbf{p}_{j}-\mathbf{p}_{i}\|\leq r,j\neq i\}$ (8)
Each deleted point’s $\epsilon_{g}$ are updated through Bayesian inference between previous $\epsilon_{g}$ and $\gamma$, similar to (7):
$\epsilon_{\mathrm{g_{t}}}=\frac{\epsilon_{\mathrm{g_{t-1}}}\cdot\gamma}{\epsilon_{\mathrm{g_{t-1}}}\cdot\gamma+(1-\epsilon_{\mathrm{g_{t-1}}})\cdot(1-\gamma)}.$ (9)
Meanwhile, emerged points $\mathcal{E}_{t}$ comprise both newly built structures and ephemeral objects, introducing uncertainty of their permanence. We thus scale their local ephemerality by an uncertainty factor $k$ along with the objectness factor:
@@
$\epsilon_{\mathrm{g,i}}=k\cdot\bigl{(}2-\gamma\bigr{)}\cdot\epsilon_{\mathrm{l,i}}\quad\forall i\in\mathcal{E}_{t}.$ (10)
@@
Points in category (iv) inherit their previous $\epsilon_{g}$ unchanged, while those in category (v) initialize $\epsilon_{g}$ directly from their $\epsilon_{l}$.
Over multiple sessions, truly static points accumulate enough observations for their $\epsilon_{g}$ to decrease and stabilize, distinguishing them from highly ephemeral objects. Changes in $\mathcal{M}_{t}$ are recorded in a delta map for session $\mathcal{S}_{t}$, which continuously tracks and logs changes ($\Delta\epsilon_{g}$). Unlike LT-mapper’s diff map [11], this delta map captures the magnitude of change, enabling fine-grained analysis of environmental variation. Finally, a static map can be retrieved by filtering $\mathcal{M}_{t}$ with a user-defined threshold $\tau_{g}$ on $\epsilon_{g}$.
@@
Figure 5: In our map update module, points are classified into five categories. Coexisting points ($\mathcal{C}_{t}$) are shown in grey. Deleted points ($\mathcal{D}_{t}$) are red if they truly disappeared and pink if they belong to previously visited regions only. Emerged points ($\mathcal{E}_{t}$) are blue if newly added and sky blue if observed only in the current session. Each category follows a specific update strategy for robust map maintenance.